# Data-Driven Room Acoustic Modeling Via Differentiable Feedback Delay Networks With Learnable Delay Lines

### v2

@@ Abstract
Automatic differentiation, differentiable feedback delay networks, room acoustics.
@@ 1 Introduction
Room acoustic synthesis involves simulating the acoustic response of an environment, a task that finds application in a variety of fields, e.g. in music production, to artistically enhance sound recordings; in architectural acoustics, to improve the acoustics of performance spaces; or in VR/AR/computer games, to enhance listeners’ sense of realism [1], immersion [2], and externalization [3].
@@
This paper proposes a new method for automatic FDN parameter tuning. The present work is rooted in a recent framework for the parameter estimation of lumped-element models [19] based on automatic differentiation [20], and its novelty is twofold. First, the cost function combines two objective measures of perceptual features, i.e., the energy decay curve (EDC) and a differentiable version of the normalized echo density profile (EDP) [21]. Second, the delay line lengths are optimized via backpropagation along with every other FDN parameter, thus allowing exploiting the flexibility of delay-network models to the fullest. We thus introduce a simple, robust, and fully automatic method for matching acoustic measurements at scale. The learned parameters can be then seamlessly plugged into off-the-shelf FDN software without further processing or mapping.
The paper is organized as follows. Section 2 introduces the background information on FDNs. Section 3 discusses the prior art on automatic FDN parameter tuning. Section 4 describes the proposed method, and Section 5 presents its evaluation. Section 6 concludes the manuscript.
@@ 2 Feedback Delay Networks
The block diagram of a single-input-single-output (SISO) FDN is show in Figure 1. This system is characterized by [22]
@@
where $u[n]$ is the input signal, $y[n]$ is the output signal, $\mathbf{b}\in\mathbb{R}^{N}$ is a vector of input gains, $\mathbf{c}\in\mathbb{R}^{N}$ is a vector of output gains, $(\cdot)^{T}$ denotes the transpose operation, $\mathbf{A}\in\mathbb{R}^{N\times N}$ is the feedback matrix, $d\in\mathbb{R}$ is the scalar gain associated to the direct path, and $\bm{m}=[m_{1},...,m_{N}]$ is a vector containing the length of the $N$ delay lines expressed in samples. The vector $\mathbf{s}[n]\in\mathbb{R}^{N}$ denotes the output of the delay lines at time index $n$, and we use the following notation $\mathbf{s}[n+\bm{m}]=\left[s_{1}[n+m_{1}],...,s_{N}[n+m_{N}]\right]^{T}$ to indicate $N$ parallel delay operations applied to $\mathbf{s}[n]$ of $m_{1},...,m_{N}$ samples, respectively.
@@
While it is possible to design feedback matrices as time varying and/or frequency dependent [24, 7], this paper focuses on the time-invariant and frequency-independent case. With this assumption, the stability of the system can be easily enforced throughout the training, thanks to the model reparameterization strategies discussed later in Section 4.
Moreover, such a model benefits from having a low computational complexity. From visual inspection of Figure 1, indeed, an FDN only requires $12N+$ 1 multiplications, $2N$ additions and one vector-matrix multiplication per sample. The vector-matrix multiplication requires $N^{2}$ scalar multiplications and $N(N-1)$ additions for the case of a generic feedback matrix (while it becomes $O(N)$ for a Householder matrix). Assuming equal cost of additions and multiplications, the overall computational complexity of an FDN amounts to $f_{\mathrm{s}}\>(2N^{2}+3N+1)$ floating-point operations per second (FLOPS), where $f_{\mathrm{s}}$ is the sampling rate. For $N=6$ and $f_{\mathrm{s}}=44.1$ kHz, that corresponds to a computational complexity of 4 MFLOPS. For comparison, modeling a $0.5$ seconds long RIR using an FIR filter (i.e., naive convolution) at the same sampling rate would carry a computation complexity of 1945 MFLOPS. In real-time applications one would normally use faster methods such as partitioned convolution [5] or overlap-add (FFT-based) convolution [25]. Under the same conditions and assuming a frame refresh rate of $50$ Hz, overlap-add convolution carries a complexity of 207 MFLOPS  [10], which is still nearly two orders of magnitude larger than an FDN.
@@ 3 Related Work
As mentioned earlier, automatic tuning of FDN parameters has been previously investigated by means of gradient-free methods, such as Bayesian optimization [12] and genetic algorithms [13, 15, 14, 16], as well as gradient-based machine learning techniques [17, 18].
Some works are concerned with the automatic tuning of off-the-shelf reverberation plug-ins. In [26], Heise and colleagues investigate four gradient-free optimization strategies: simulated evolution [27], the Nelder-Mead simplex method [28], Nelder-Mead with brute-force parallelization, and particle swarm optimization [29]. More recently, [12] applies Bayesian optimization using a Gaussian process as a prior to iteratively acquire the control parameters of an external FDN plug-in that minimize the mean absolute error between the multiresolution mel-spectrogram of the target RIR convolved with a 3-second logarithmic sine sweep and that of the artificial reverberator output. The FDN control parameters include the delay line length, reverberation time, fade-in time, high/low cutoff, high/low Q, high/low gain, and dry-wet ratio.
Conversely, other studies assume to have white-box access to the delay-network structure and apply genetic algorithms (GA) to optimize a subset of the FDN parameters. In [13], a GA is used to find both the $N^{2}$ coefficients of the feedback matrix $\mathbf{A}$ and $N$ cutoff frequencies of lowpass filters, one for each delay line. The authors of [14] aim at finding a mapping between room and FDN parameters for VR/AR applications. To this end, they synthesize the binaural RIR of a set of virtual shoebox rooms, apply a GA to tune the FDN’s delay lines and scalar feedback gain, and use the resulting training pairs to fit a support vector machine (SVM) regressor. In [15], Coggin and Pirkle apply a GA for the estimation of $\bm{m}$, $\mathbf{b}$, and $\mathbf{c}$. For every individual in a generation, attenuation and output filters are designed using the Yule-Walker method. The authors investigate several fitness functions before favoring the Chebyshev distance between the power envelopes of the target and predicted IR. The optimization is run for late reverberation only: the first 85 ms of the RIR are cut and convolved with the input signal, before being fed to the FDN to model late reverberation. Following [15], Ibnyahya and Reiss [16] recently introduced a multi-stage method combining more advanced analytical filter design methods and GAs to estimate the FDN parameters that would best approximate a target RIR in terms of a MFCC-based fitness function, similarly to the cost function used in [26].
@@
In a different vein, several recent works aim at learning lumped parameters via gradient-based optimization directly within the digital structure of the model, and forgo parameter-yielding neural networks altogether. In this respect, automatic differentiation has been recently proposed to find $\mathbf{A},\mathbf{b},\mathbf{c}$ of an FDN (without parameterizing them as a neural network) so as to minimize spectral coloration and obtain a flat frequency response [18]. Similar yet distinct, other works adopt a white-box system identification approach and use backpropagation to find the parameters of predetermined mathematical models so as to match measured data as closely as possible [32, 33, 19].
In this work, we adopt the latter approach and use the method detailed in the next section to find the values of $\mathbf{A}$, $\mathbf{b}$, $\mathbf{c}$, $\bm{m}$, and $d$ such that the resulting FDN is capable of modeling perceptually meaningful characteristics of the acoustic response of real-life environments
@@ 4 Proposed Method
At training time, we strip out the initial silence due to direct-path propagation, and disregard every sample beyond the $T_{60}$ of the target RIR. In other words, we only consider the first $L_{T_{60}}:=\left\lceil T_{60}\cdot f_{\mathrm{s}}\right\rceil$ samples of $h[n]$ and $\hat{h}[n]$ in computing the loss. The reason behind restricting the temporal scope only to the segment of the RIR associated with the $T_{60}$ is that, beyond this point, the values involved in the ensuing computations become so small that numerical errors might occur when using single-precision floating-point numbers, and the training process would unwantedly focus on the statistics of background/numerical noise. Notice that, at inference time, i.e., once the FDN parameters have been learned, the room acoustics simulation can be run indefinitely at a very low computational cost.
@@ 4.2 Trainable Gains
We would like the input, output, and direct gains of our differentiable FDN to be nonnegative. This way, the gains only affect the amplitude of the signals and do not risk inverting their polarity. Instead, we let $\mathbf{A}$ model phase-reversing reflections. To enforce gain nonnegativity, we employ a differentiable nonlinear function $f_{\geq 0}:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$, such as, e.g., the Softplus or exponential function. We then learn, e.g., $\mathbf{\tilde{b}}=[\tilde{b}_{1},...,\tilde{b}_{N}]^{T}$ while using $\mathbf{b}=[f_{\geq 0}(\tilde{b}_{1}),...,f_{\geq 0}(\tilde{b}_{N})]^{T}$ in every computation concerning the FDN. Among other options, we select $f_{\geq 0}(x)=\lvert x\rvert$ [19], where the requirement of $f_{\geq 0}$ being differentiable everywhere was relaxed as it is common for many widely-adopted activation functions, such as ReLU.
@@ 4.3 Trainable Feedback Matrix
We focus on lossy FDNs. In prior work [18], frequency-independent homogeneous decay has been modeled by parameterizing $\mathbf{A}$ as the product of a unilossless matrix $\mathbf{U}$ and a diagonal matrix $\bm{\Gamma}(\bm{m})=\operatorname{diag}\left(\gamma_{1},...,\gamma_{N}\right)=\operatorname{diag}\left(\gamma^{m_{1}},...,\gamma^{m_{N}}\right)$ containing a delay-dependent absorption coefficient for each delay line, where $\gamma\in(0,1)$ is a constant gain-per-sample coefficient. The feedback matrix is thus expressed as
@@
As for the matrix $\bm{\Gamma}(\bm{m})$, we noticed that tying the values of the absorption coefficients $\gamma_{1},...,\gamma_{N}$ to those of the fractional delays $m_{1},...,m_{N}$ led to instability during training since the values in $\bm{m}$ were concurrently acting on the temporal location of the IR taps as well as their amplitude. [^2] Notice that this problem is unique to our approach, as previous studies employed non-trainable delay lines with fixed lengths [18]. Conversely, we decouple $\bm{\Gamma}$ from $\bm{m}$, thus learning a possibly inhomogeneous FDN characterized by $\mathbf{A}=\mathbf{U}\bm{\Gamma}$, as opposed to the homogeneous FDNs studied in [18].
@@ 4.4 Trainable Delay Lines
In the digital domain, an integer delay can be efficiently implemented as a reading operation from a buffer that accumulates past samples. Unfortunately, this approach is non-differentiable. Instead, since the Fourier transform is a linear and differentiable operator, we may think to work in the frequency domain to circumvent such a problem. In particular, we evaluate the delays on the unit circle at discrete frequency points $\omega_{k}=2\pi k/K$, for $k=0,...,K-1$.
Each differentiable delay line embeds a $K$-sample buffer. First, we compute the Fast Fourier Transform (FFT) of the signal $x[n]$ currently stored in the buffer. Then, we apply a (fractional) delay of $m$ samples by multiplying the resulting spectrum with the complex exponential $e^{-j\omega_{k}m}$. Finally, we go back in the time domain by computing the inverse FFT. We can express this sequence of differentiable operations as
$x[n-m]=\frac{1}{K}\sum_{k=0}^{K-1}\underbrace{\left(\sum_{\tau=0}^{K-1}x[\tau]e^{-j\omega_{k}\tau}\right)}_{\operatorname{FFT}\left\{x[n]\right\}}e^{j\omega_{k}(n-m)}$ (5)
Optionally, cosine-tapered windows can be included in (5) to reduce spectral leakage. Nevertheless, preliminary experiments involving Hann windowing showed no meaningful improvement in the method’s performance.
Delays must be nonnegative to realize a casual system. Hence, we use $f_{\geq 0}$ to ensure nonnegativity of the fractional delays, i.e., $m_{i}=f_{\geq 0}(\tilde{m}_{i})$ where $\tilde{m}_{i}\in\mathbb{R}$ is the $i$th trainable delay-line length proxy, $i=1,...,N$. Notice that such a parameterization, while enforcing nonegativity, imposes no upper bound on $m_{i}$. An alternative parameterization is $m_{i}=K\cdot f_{(0,1)}(\tilde{m}_{i})$, which ensures that $m_{i}\in(0,K)$ at all times. In this work, we favored the former for its lower computational cost.
Finally, it is worth highlighting three main reasons for labeling our FDN model as time-domain, despite the implementation of differentiable delay lines in the frequency domain. First, frequency-domain operations are confined within the delay filterbank. Second, we stress that our FDN yields the output one sample at a time according to (1). Third, we emphasize the difference between our approach and existing methods implementing every FDN operation in the frequency domain [17, 18].

### v3

@@ Abstract
Automatic differentiation, feedback delay networks, room acoustics.
Peer-Reviewed Version
A peer-reviewed version of the manuscript has been published in EURASIP Journal of Audio, Speech, and Music Processing and is available at 10.1186/s13636-024-00371-5. Please, cite this work as: A. I. Mezza, R. Giampiccolo, E. De Sena, and A. Bernardini, ”Data-Driven Room Acoustic Modeling Via Differentiable Feedback Delay Networks With Learnable Delay Lines,” EURASIP Journal of Audio, Speech, and Music Processing, vol. X, pp. X-X, 2024, doi: 10.1186/s13636-024-00371-5.
@article{mezza2024dfdn,
title = {Data-Driven Room Acoustic Modeling Via Differentiable Feedback Delay Networks With Learnable Delay Lines},
author = {Alessandro Ilic Mezza and Riccardo Giampiccolo and Enzo De Sena and Alberto Bernardini},
journal = {EURASIP Journal of Audio, Speech, and Music Processing},
volume = {X},
pages = {X-X},
year = {2024},
issn = {1687-4722},
doi = {https://doi.org/10.1186/s13636-024-00371-5}
@@ 1 Introduction
Room acoustic synthesis involves simulating the acoustic response of an environment, a task that finds application in a variety of fields, e.g., in music production, to artistically enhance sound recordings; in architectural acoustics, to improve the acoustics of performance spaces; or in VR/AR/computer games, to enhance listeners’ sense of realism [1], immersion [2], and externalization [3].
@@
This paper proposes a new method for automatic FDN parameter tuning. The present work is rooted in a recent framework for the parameter estimation of lumped-element models [19] based on automatic differentiation [20], and its novelty is twofold. First, the cost function combines two objective measures of perceptual features, i.e., the Energy Decay Curve (EDC) and a differentiable version of the normalized Echo Density Profile (EDP) [21]. Second, the delay line lengths are optimized via backpropagation along with every other FDN parameter, thus allowing exploiting the flexibility of delay-network models to the fullest. We thus introduce a simple, robust, and fully automatic method for matching acoustic measurements. The learned parameters can be then seamlessly plugged into off-the-shelf FDN software without further processing or mapping.
The paper is organized as follows. Section 2 introduces the background information on FDNs. Section 3 discusses the prior art on automatic FDN parameter tuning. Section 4 describes the proposed method, and Section 5 presents its evaluation. Finally, Section 6 concludes the manuscript.
@@ 2 Feedback Delay Networks
The block diagram of a single-input-single-output (SISO) FDN is shown in Figure 1. This system is characterized by [22]
@@
where $u[n]$ is the input signal, $y[n]$ is the output signal, $\mathbf{b}\in\mathbb{R}^{N}$ is a vector of input gains, $\mathbf{c}\in\mathbb{R}^{N}$ is a vector of output gains, $(\cdot)^{T}$ denotes the transpose operation, $\mathbf{A}\in\mathbb{R}^{N\times N}$ is the feedback matrix, $d\in\mathbb{R}$ is the scalar gain associated to the direct path, and $\bm{m}=[m_{1},...,m_{N}]$ is a vector containing the length of the $N$ delay lines expressed in samples. The vector $\mathbf{s}[n]\in\mathbb{R}^{N}$ denotes the output of the delay lines at time index $n$, and we use the following notation $\mathbf{s}[n+\bm{m}]=\left[s_{1}[n+m_{1}],...,s_{N}[n+m_{N}]\right]^{T}$ to indicate $N$ parallel delay operations of $m_{1},...,m_{N}$ samples, respectively, applied to $\mathbf{s}[n]$.
@@
While it is possible to design feedback matrices as time varying and/or frequency dependent [24, 7], this paper focuses on the time-invariant and frequency-independent case. With this assumption, the stability of the system can be easily enforced throughout the training, thanks to the model reparameterization strategies discussed later in Section 4. Moreover, time-invariant frequency-independent FDNs benefit from having a low computational complexity. From visual inspection of Figure 1, indeed, such an FDN only requires $12N+$ 1 multiplications, $2N$ additions and one vector-matrix multiplication per sample. The vector-matrix multiplication requires $N^{2}$ scalar multiplications and $N(N-1)$ additions for the case of a generic feedback matrix (while it becomes $O(N)$ for a Householder matrix). Assuming equal cost of additions and multiplications, the overall computational complexity of an FDN amounts to $f_{\mathrm{s}}\>(2N^{2}+3N+1)$ floating-point operations per second (FLOPS), where $f_{\mathrm{s}}$ is the sampling rate. For $N=6$ and $f_{\mathrm{s}}=44.1$ kHz, that corresponds to a computational complexity of $4$ MFLOPS. For comparison, modeling a $0.5$ s long RIR using an FIR filter (i.e., naive convolution) at the same sampling rate would carry a computation complexity of $1945$ MFLOPS. In real-time applications, one would normally use faster methods such as partitioned convolution [5] or overlap-add (FFT-based) convolution [25]. Under the same conditions and assuming a frame refresh rate of $50$ Hz, overlap-add convolution carries a complexity of $207$ MFLOPS  [10], which is still nearly two orders of magnitude larger than an FDN.
@@ 3 Related Work
As mentioned earlier, the automatic tuning of FDN parameters has been previously investigated by means of gradient-free methods, such as Bayesian optimization [12] and genetic algorithms [13, 15, 14, 16], as well as gradient-based machine learning techniques [17, 18].
Some works are concerned with the automatic tuning of off-the-shelf reverberation plug-ins. In [26], Heise and colleagues investigate four gradient-free optimization strategies: simulated evolution [27], the Nelder-Mead simplex method [28], Nelder-Mead with brute-force parallelization, and particle swarm optimization [29]. More recently, [12] applies Bayesian optimization using a Gaussian process as a prior to iteratively acquire the control parameters of an external FDN plug-in that minimize the mean absolute error between the multiresolution mel-spectrogram of the target RIR convolved with a 3 s logarithmic sine sweep and that of the artificial reverberator output. The FDN control parameters include the delay line length, reverberation time, fade-in time, high/low cutoff, high/low Q, high/low gain, and dry-wet ratio.
Conversely, other studies assume to have white-box access to the delay-network structure and apply genetic algorithms (GA) to optimize a subset of the FDN parameters. In [13], a GA is used to find both the $N^{2}$ coefficients of the feedback matrix $\mathbf{A}$ and $N$ cutoff frequencies of lowpass filters, one for each delay line. The authors of [14] aim at finding a mapping between room and FDN parameters for VR/AR applications. To this end, they synthesize the binaural RIRs of a set of virtual shoebox rooms, apply a GA to tune the FDN’s delay lines and scalar feedback gain, and use the resulting training pairs to fit a support vector machine (SVM) regressor. In [15], Coggin and Pirkle apply a GA for the estimation of $\bm{m}$, $\mathbf{b}$, and $\mathbf{c}$. For every individual in a generation, attenuation and output filters are designed using the Yule-Walker method. The authors investigate several fitness functions before favoring the Chebyshev distance between the power envelopes of the target and predicted IR. The optimization is run for late reverberation only: the first $85$ ms of the RIR are cut and convolved with the input signal, before being fed to the FDN to model late reverberation. Following [15], Ibnyahya and Reiss recently introduced a multi-stage method [16] combining more advanced analytical filter design methods and GAs to estimate the FDN parameters that would best approximate a target RIR in terms of an MFCC-based fitness function similar to the cost function used in [26].
@@
In a different vein, several recent works aim at learning lumped parameters via gradient-based optimization directly within the digital structure of the model and forgo parameter-yielding neural networks altogether. In this respect, automatic differentiation has been recently proposed to find $\mathbf{A},\mathbf{b},and\mathbf{c}$ of an FDN (without parameterizing them as a neural network) so as to minimize spectral coloration and obtain a flat frequency response [18]. Similar yet distinct, other works adopt a white-box system identification approach and use backpropagation to find the parameters of predetermined mathematical models so as to match measured data as closely as possible [32, 33, 19].
In this work, we adopt the latter approach and use the method detailed in the next section to find the values of $\mathbf{A}$, $\mathbf{b}$, $\mathbf{c}$, $\bm{m}$, and $d$ such that the resulting FDN is capable of modeling perceptually meaningful characteristics of the acoustic response of real-life environments.
@@ 4 Proposed Method
At training time, we strip out the initial silence due to direct-path propagation and disregard every sample beyond the $T_{60}$ of the target RIR. In other words, we only consider the first $L_{T_{60}}:=\left\lceil T_{60}\cdot f_{\mathrm{s}}\right\rceil$ samples of $h[n]$ and $\hat{h}[n]$ in computing the loss. The reason behind restricting the temporal scope only to the segment of the RIR associated with the $T_{60}$ is that, beyond this point, the values involved in the ensuing computations become so small that numerical errors might occur when using single-precision floating-point numbers, and the training process would unwantedly focus on the statistics of background/numerical noise. Notice that, at inference time, i.e., once the FDN parameters have been learned, the room acoustics simulation can be run indefinitely at a very low computational cost.
@@ 4.1 Model Reparameterization
Figure 2 summarizes the proposed approach, listing all the unconstrained trainable parameters (top), the corresponding reparameterization (middle), and illustrating once again how, through (1), the FDN processes time-domain signals, including unit impulses $\delta[n]$ (bottom).
Figure 2: Summary of the proposed method.
@@ 4.2 Trainable Gains
We would like the input, output, and direct gains of our differentiable FDN to be nonnegative. This way, the gains only affect the amplitude of the signals and do not risk inverting their polarity. Instead, we let $\mathbf{A}$ model phase-reversing reflections. To enforce gain nonnegativity, we employ a differentiable nonlinear function $f_{\geq 0}:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$, such as the Softplus or exponential function. We then learn, e.g., $\mathbf{\tilde{b}}=[\tilde{b}_{1},...,\tilde{b}_{N}]^{T}$ while using $\mathbf{b}=[f_{\geq 0}(\tilde{b}_{1}),...,f_{\geq 0}(\tilde{b}_{N})]^{T}$ in every computation concerning the FDN. Among other options, we select $f_{\geq 0}(x)=\lvert x\rvert$ [19], where the requirement of $f_{\geq 0}$ being differentiable everywhere was relaxed as it is common for many widely-adopted activation functions, such as ReLU.
@@ 4.3 Trainable Feedback Matrix
We focus on lossy FDNs. In prior work [18], frequency-independent homogeneous decay has been modeled by parameterizing $\mathbf{A}$ as the product of a unilossless matrix $\mathbf{U}$ and a diagonal matrix $\bm{\Gamma}(\bm{m})=\operatorname{diag}\left(\gamma_{1},...,\gamma_{N}\right)=\operatorname{diag}\left(\gamma^{m_{1}},...,\gamma^{m_{N}}\right)$ containing a delay-dependent absorption coefficient for each delay line, where $\gamma\in(0,1)$ is a constant gain-per-sample parameter. The feedback matrix is thus expressed as
@@
As for the matrix $\bm{\Gamma}(\bm{m})$, we noticed that tying the values of the absorption coefficients $\gamma_{1},...,\gamma_{N}$ to those of the fractional delays $m_{1},...,m_{N}$ as previously done in [18] led to instability during training since the values in $\bm{m}$ were concurrently acting on the temporal location of the IR taps as well as their amplitude. [^2] This problem is unique to our approach, as previous studies employed non-trainable delay lines with fixed lengths [18]. Conversely, we decouple $\bm{\Gamma}$ from $\bm{m}$, thus learning a possibly inhomogeneous FDN characterized by $\mathbf{A}=\mathbf{U}\bm{\Gamma}$, as opposed to the homogeneous FDNs studied in [18].
@@ 4.4 Trainable Delay Lines
In the digital domain, an integer delay can be efficiently implemented as a reading operation from a buffer that accumulates past samples. Unfortunately, this approach is not differentiable. Instead, since the Fourier transform is a linear and differentiable operator, we opt to work in the frequency domain to circumvent the problem.
In [36], Pei and Lai proposed a closed-form variable fractional delay filter, which turns out to be inherently differentiable. In our FDN implementation, each delay line is equipped with a $Q$-sample buffer, so that the $i$th buffer stores the signal $x_{i}[n]$. First, we zero-pad $x_{i}[n]$ to reduce artifacts due to the ensuing circular convolution. Then, we compute the $K$-point Fast Fourier Transform (FFT) of the resulting signal, with $K=2Q$. Following [36], we apply a delay of $m_{i}$ (fractional) samples by multiplying the discrete spectrum with the conjugate symmetric frequency response $D_{i}[k]$ defined in (6). Finally, we go back in the time domain by computing the inverse FFT. We can express this sequence of differentiable operations as
$x_{i}[n-m]=\operatorname{IFFT}\left\{D_{i}[k]\cdot\operatorname{FFT}\left\{x_{i}[n]\right\}\right\},$ (5)
where
$D_{i}[k]=\begin{cases}1,&k=0\\ e^{-jm_{i}(2\pi/K)k},&k=1,...,\frac{K}{2}-1\\ \cos(m_{i}\pi),&k=\frac{K}{2}\\ e^{-jm_{i}(2\pi/K)(K-k)},&k=\frac{K}{2}+1,...,K-1\\ \end{cases}$ (6)
which, in the time domain, corresponds to a windowed-sinc finite impulse response [36].
Delays $m_{1},...,m_{N}$ must be nonnegative to realize a casual system. Hence, we use $f_{\geq 0}$ to reparameterize them. Moreover, we clip the resulting values so not to exceed the given buffer length. This yields333 It is worth pointing out that (7) is not the only way to account for the implicit periodization of the buffered signal when computing the FFT. For instance, an alternative parameterization is $m_{i}=(Q-1)\cdot f_{(0,1)}(\tilde{m}_{i})$, which ensures that $m_{i}\in(0,Q-1)$ at all times.
$m_{i}=\min\left(Q-1,f_{\geq 0}(\tilde{m}_{i})\right),$ (7)
where $\tilde{m}_{i}\in\mathbb{R}$ is the $i$th trainable delay-line length proxy, $i=1,...,N$.
Finally, it is worth highlighting three main reasons for labeling our FDN model as time-domain, despite implementing the differentiable delay lines in the frequency domain. First, we stress that our FDN yields the output one sample at a time according to (1). Second, frequency-domain operations are confined within the delay filterbank. Since delay lines being differentiable is only required at training time, the inference model can thus feature a different fractional delay implementation, possibly in the time domain. Third, we emphasize the difference between our approach and existing methods implementing every FDN operation in the frequency domain [17, 18].